OpenStudy (anonymous):

You deposit $10,000 in an account earning 4% interest compounded monthly. a. How much will you have in the account in 25 years? b. How much interest will you earn?

OpenStudy (anonymous):

a. P = Po(1+r)^n. Po = $10,000. r = (4%/12) / 100% = 0.0033333 = Monthly % rate expressed as a decimal. n = 12Comp./yr * 25yrs = 300 Compounding periods. Plug the above values into the given Eq. Answer: P = $27,137.65. Im not 100% sure on this but I think this is right